# Cybersecurity in Critical Infrastructures: A Post-Quantum Cryptography Perspective

### v1

@@ Abstract
The machinery of industrial environments was connected to the Internet years ago with the scope of increasing their performance. However, this made such environments vulnerable against cyber-attacks that can compromise their correct functioning resulting in economic or social problems. Lately, an increase of cyberattacks to industrial environments has been experienced as a result of their lack of cybersecurity resources. Moreover, implementing cryptosystems in the communications between Operational Technologies (OT) devices is a more challenging task than for Information Technologies (IT) environments since the OT are generally composed of legacy elements, characterized by low-computational capabilities or non-standardized communication protocols. Consequently, implementing cryptosystems in industrial communication networks faces a trade-off between the security of the communications and the amortization of the industrial infrastructure. Critical Infrastructure (CI) refers to the industries which provide key resources for the daily social and economical development, e.g. electricity or water. CI communications are a very exposed target to cyberattacks that could produce huge economical and social losses, even human casualties. Furthermore, a new threat to cybersecurity has arisen with the theoretical proposal of fault tolerant quantum computers, due to their potential ability of breaking state-of-the-art cryptography protocols, such as Rivest-Shamir-Adleman (RSA) or Elliptic-Curve Cryptography (ECC). The chase of functional quantum computers has resulted in a technological race involving many global agents. At the same time, those agents have become aware that transitioning their secure communications to a quantum secure paradigm is a priority that should be established before the arrival of fault-tolerance. In this sense, two main cryptographic solutions have been proposed: Quantum key Distribution (QKD) and Post-Quantum Cryptography (PQC). Nevertheless, quantum secure solutions have been mainly centered from the perspective of IT environments. In this paper, we provide a perspective of the problem of applying PQC solutions to CI and analyze which could be the most suitable cryptography schemes for these scenarios.
@@ I Introduction
[…]
@@
In this sense, relying on communications for executing the critical tasks involved in such hyperconnected paradigm requires that those transmissions of information are secure and private. Cyber vulnerabilities in the control systems of a smart city or an automated industry may lead to catastrophic consequences. For example, in a possible future where the transport of people and cargo is exclusively done by autonomous vehicles which rely on the communications among them and some central control stations to move around, the intrusion of a malicious entity on the system to disturb it would lead to fatal consequences economically and socially (casualties). Hence, modern crime and war is heavily based on hacking activities with the scope of manipulating critical infrastructures (CI), to produce economical or social losses by interrupting their production or by decreasing the life-time of their devices, or obtaining sensitive information regarding state, industrial or personal secrets (banking information or sensitive images, for example). This paradigm of cybercrime and cyberwar is present nowadays with an estimated amount of $2200$ known cyberattacks per day in 2022, posing a threat to the business’ infrastructure every $39$ seconds [2]. Notoriously, the recent Russo-Ukrainian war has seen a major cyberwarfare in which the public energy, media, financial, business and non-profit sectors of Ukraine have been targeted by hacking activities that have undermined the distribution of medicines, food and relief supplies among other harmful effects [3]. Indeed, awareness on cyberattacks among Nation-state actors is increasing due to current geopolitical tensions, as seen recently [4]. It is due to all these factors that concepts such as the Cyber Apocalypse are being coined to describe the fear that a cyberattack to CI’s systems and networks of a country would led to shutting down their capabilities regarding civilian and military services. It is important to state that the possibility of major devastation in the CI of a nation does not have to imply that all the systems consisting it should be attacked, the failure of parts of the structure may led to a catastrophic propagation of failures through the whole network due to the interconnection among the elements. This effect is known as cascading effect [5].
@@
All of these vulnerabilities make cybersecurity and cryptography to be the pillars to erect the previously described paradigmatic society in a security way. Cybersecurity is defined as the practice of protecting important systems and confidential information from cyberattacks. In this sense, many methods and elements are used for the sake of protecting communication and computer networks, but the algorithms that are employed to cipher sensitive data being communicated in such meshes relate to the field of cryptography. Hence, in this society where the quote “Information is power” is getting more and more relevant, the use of such practices is of insurmountable importance. Importantly, the proposal of the RSA or ECC asymmetric cryptographic systems has maintained the security of communication systems for over 40 years [6, 7, 8]. The core of those protocols resides in the fact that they are based on hard problems that cannot be solved in a practical time frame by classical computing methods, i.e. thousands of years of computing are required to extract the plain text from the ciphertext if the key is unknown. Unluckily, quantum computers have posed a threat to the security of those asymmetric cryptography protocols. The Shor’s algorithm is a theoretical quantum algorithm that provides an exponential speedup for solving prime number factorization and the computation of discrete logarithms, which are the hard problems in which the security of the previously commented protocols rely on, respectively [9]. Quantum computers that can execute such algorithm efficiently and correctly still exist just as a theoretical promise, but the last years have shown that it is a rapidly evolving field with the achievement of milestones such as the first experimental realizations of quantum advantage [10, 11, 12, 13, 14] or quantum error correction [15, 16]. Therefore, it is a matter of time, usually estimated to be between one and two decades from now [17], that quantum computers make the state-of-the-art asymmetric protocols to be deprecated. Hence, many have raised the alarm of a possible “Quantum Apocalypse” that would result in sensitive data and systems to become completely vulnerable.
@@
Each PQC protocol has its own benefits and disadvantages in terms of security levels, ciphertext size or speed, among other benchmarks. This implies that the selection of PQC protocols is very application-dependent in the sense that as a function of the requirements of a specific system, an approach could be valid or not. Following this logic, PQC protocols are usually proposed for systems in which the cybersecurity is the most critical requirement (IT services), while the latency because of the introduction of those cryptography protocols can be deemed as not too important. However, latency is a key performance parameter in Industrial Control Systems (ICS) and CI, where introducing a delay over the system requirements can imply a failure that cannot be tolerated in such environments [28]. This should obviously be done maintaining a certain level of security on the system. Additionally, it is important to state that implementing cryptography in such networks is done by means of processors that are not powerful enough to manage huge key sizes, mainly because the introduction of such systems should be somehow seamless to the existing communication infrastructure and cheap [^1] We can speculate that QKD will not be a major player to secure such networks due the fact that they are very costly as well as they require significant infrastructure to be deployed.. It is in this sense that, the inclusion of PQC in industrial and critical environments poses an interesting trade-off between the benchmarks of those protocols. As mentioned previously, protecting ICS and CI from possible cyberattacks is fundamental due to the humongous impact that those systems have in society and industry, making their failure to cause intolerable economic losses and, in the worse scenario, injuries and even casualties. Interestingly, this systems have shown to be vulnerable in the recent times with several hacking proposals [29, 30, 31, 32]. Therefore, the necessity of transitioning the security of industrial and CI to post-quantum cryptography is central to keep all those systems secure against a possible quantum threat, as it has been recently noted by the United States Cybersecurity and Infrastructure Security Agency (CISA), the National Security Agency (NSA) and the NIST [33]. In this perspective, we review and discuss the current affairs regarding the integration of PQC security in industrial environments and CIs. Thus, in section II we provide a review of the communication systems in such environment as well as of the stringent requirements for integrating cybersecurity in them. We follow, in section III, with a short review on cryptography and the threat that quantum computers pose to traditional ciphering schemes. Existing proposals of PQC algorithms are surveyed in section IV presenting several protocols proposed by many worldwide agents. In addition, the performance benchmarks of those PQC candidates are presented. An overview of the state of PQC in industrial and CI environments is finally presented in section V, were we speculatively discuss which existing protocols may be the ones for integration in those scenarios.
@@ II Industrial environments
ICS are the components of the industrial sector and infrastructures, from essential services such as energy, water, transportation systems to manufacturing plants, agricultural systems, building automation systems, etc. In these infrastructures we will always find complex components that share a common denominator: physical processes that are modified by logical computation or viceversa. These components are called Cyber-Physical Systems (CPS). In figure 1, we show an example of ICS. In this example, in short, the instrumentation sensors measure physical variables, programmable logic controllers (PLCs) implement a control loop and send signals to the actuators. All the process is controlled and monitored by the Supervisory Control And Data Acquisition (SCADA) and the operator interacts via the Human-Machine Interface (HMI) or the Engineering Work Station (EWS).
@@
Since these systems control real world processes, any potential cyberattack impact on them could imply a physical effect in the real world. Hence, cyber-risks could turn not only into production downtimes but also physical damage to operators or users.
ICS and CPS are a part of larger infrastructures which interact with information technology (IT) systems at certain point. In figure 2, we show how IT and operational technology (OT) interact in this type of infrastructures according to the Purdue model, even though there are several models such as RIA 4.0 and others.
@@ II.1 Differences between IT and OT
In summary, the cost of achieving a high level of cybersecurity can be much higher in older industrial plants due to the need for retrofitting and upgrading legacy systems. In contrast, new companies and modern facilities have the advantage of incorporating cybersecurity measures at a lower cost from the outset, thanks to the availability of cyber-embedded components and systems. However, it is crucial for all organizations, regardless of age, to prioritize cybersecurity to protect critical infrastructure and assets from evolving cyber threats.
II.2 Industrial Cybersecurity Standards and Mechanisms
@@
In addition to broader standards and regulations, there are specific standards for components and systems used in industrial environments. These may include industrial communication protocols and HW requirements. By adhering to these standards, organizations can ensure compatibility and security among different components and systems within their infrastructure. The interoperability provided by these specific standards is crucial for maintaining a secure and efficient industrial ecosystem. Figure 3 summarizes these interdependencies.
@@ III.1 Overview on cryptography
In the following, we provide a comprehensive explanation of both schemes based on a network composed of three parties: a server, a user who wants to have a secure connection with it and an eavesdropper who wants to obtain information about the message. In cryptography those are usually referred to as Alice, Bob and Eve, respectively.
III.1.1 Symmetric Key Schemes
@@
III.1.2 Asymmetric Key Schemes
As defined above, asymmetric key or public key schemes are defined as cryptosystems where each of the parties involved in the communication use their own key to secure the information, the private key and the public key. Each of those are usually employed to share a secret between two parts and, therefore, this encryption schemes may be used for the confidential key exchange required in symmetric schemes. In this sense, this could refer to the transmission of the bit string used as the input of an algorithm which can be used by Alice and Bob to generate the same symmetric key. In asymmetric key schemes confidentiality, integrity and authenticity are achieved in the following ways:

### v2

@@ Abstract
The machinery of industrial environments was connected to the Internet years ago with the scope of increasing their performance. However, this change made such environments vulnerable against cyber-attacks that can compromise their correct functioning resulting in economic or social problems. Moreover, implementing cryptosystems in the communications between operational technology (OT) devices is a more challenging task than for information technology (IT) environments since the OT networks are generally composed of legacy elements, characterized by low-computational capabilities. Consequently, implementing cryptosystems in industrial communication networks faces a trade-off between the security of the communications and the amortization of the industrial infrastructure. Critical Infrastructure (CI) refers to the industries which provide key resources for the daily social and economical development, e.g. electricity. Furthermore, a new threat to cybersecurity has arisen with the theoretical proposal of quantum computers, due to their potential ability of breaking state-of-the-art cryptography protocols, such as RSA or ECC. Many global agents have become aware that transitioning their secure communications to a quantum secure paradigm is a priority that should be established before the arrival of fault-tolerance. In this paper, we aim to describe the problematic of implementing post-quantum cryptography (PQC) to CI environments. For doing so, we describe the requirements for these scenarios and how they differ against IT. We also introduce classical cryptography and how quantum computers pose a threat to such security protocols. Furthermore, we introduce state-of-the-art proposals of PQC protocols and present their characteristics. We conclude by discussing the problematic of integrating PQC in industrial environments.
@@ I Introduction
The exponential development of communication technologies in the late 20th century and, specially, in the early 21st century has resulted in a contemporary society that exists in a hyperconnected world. In this paradigm, communications do not only refer to the actions of texting, phone (video) calls, social media or news media but also to the control of industrial machines, bank transfers, stock acquisitions, control of unmanned aerial vehicles (UAV) or managing automated houses (domotics), to name a few. Furthermore, strongly tangled concepts such as Smart Cities, Industry 4.0 or the Internet of Things (IoT) are currently being investigated for their convergence with other advanced technologies such as Artificial Intelligence (AI) or Quantum Computing (QC) on a historical inflection point in the form of a Fourth Industrial Revolution [1] .
@@
In this sense, relying on communications for executing the critical tasks involved in such hyperconnected paradigm requires that those transmissions of information are secure and private. Cyber vulnerabilities in the control systems of a smart city or an automated industry may lead to catastrophic consequences. For example, in a possible future where the transport of people and cargo is exclusively done by autonomous vehicles which rely on the communications among them and some central control stations to move around, the intrusion of a malicious entity on the system to disturb it would lead to fatal consequences economically and socially (casualties). Hence, modern crime and war is heavily based on hacking activities with the scope of manipulating critical infrastructures (CI), to produce economical or social losses by interrupting their production or by decreasing the life-time of their devices, or obtaining sensitive information regarding state, industrial or personal secrets (banking information or sensitive images, for example). This paradigm of cybercrime and cyberwar is present nowadays with an estimated amount of $2200$ known cyberattacks per day in 2022, posing a threat to the business’ infrastructure every $39$ seconds [2]. Indeed, awareness on cyberattacks among Nation-state actors is increasing due to current geopolitical tensions, as seen recently [3]. It is due to all these factors that concepts such as the Cyber Apocalypse are being coined to describe the fear that a cyberattack to CI’s systems and networks of a country would led to shutting down their capabilities regarding civilian and military services. It is important to state that the possibility of major devastation in the CI of a nation does not have to imply that all the systems consisting it should be attacked, the failure of parts of the structure may lead to a catastrophic propagation of failures through the whole network due to the interconnection among the elements. This effect is known as cascading effect [4].
@@
All of these vulnerabilities make cybersecurity and cryptography to be the pillars to erect the previously described paradigmatic society in a security way. Cybersecurity is defined as the practice of protecting important systems and confidential information from cyberattacks. In this sense, many methods and elements are used for the sake of protecting communication and computer networks, but the algorithms that are employed to cipher sensitive data being communicated in such meshes relate to the field of cryptography. Hence, in this society where the quote “Information is power” is getting more and more relevant, the use of such practices is of capital relevance. Importantly, the proposal of the RSA or ECC asymmetric cryptographic systems has maintained the security of communication systems for over 40 years [5, 6, 7]. The core of those protocols resides in the fact that they are based on hard problems that cannot be solved in a practical time frame by classical computing methods, i.e. thousands of years of computing are required to extract the plain text from the ciphertext if the key is unknown. Unluckily, quantum computers have posed a threat to the security of those asymmetric cryptography protocols. Shor’s algorithm is a theoretical quantum algorithm that provides an exponential speedup for solving prime number factorization and the computation of discrete logarithms respectively, which are the hard problems in which the security of the previously commented protocols relies upon [8]. At the current time, quantum computers that can execute such algorithm efficiently and correctly only exists as a theoretical promise. Nonetheless, the past years, quantum computing has proven to be a rapidly evolving field with the achievement of milestones such as the first experimental realizations of quantum advantage [9, 10, 11, 12, 13] or quantum error correction [14, 15]. Such tremendous advancements have made educated voices to estimate the appearance of efficient quantum computers able to make state-of-the-art asymmetric protocols to be deprecated to be within the range of one to two decades [16]. Hence, many have raised the alarm of a possible “Quantum Apocalypse” that would result in sensitive data and systems to become completely vulnerable.
@@
Each PQC protocol has its own benefits and disadvantages in terms of security levels, ciphertext size or speed, among other benchmarks. This implies that the selection of PQC protocols is very application-dependent in the sense that as a function of the requirements of a specific system, an approach could be valid or not. Following this logic, PQC protocols are usually proposed for systems in which the cybersecurity is the most critical requirement (IT services), while the latency because of the introduction of those cryptography protocols can be deemed as not too important. However, latency is a key performance parameter in Industrial Control Systems (ICS) and CI, where introducing a delay over the system requirements can imply a failure that cannot be tolerated in such environments [27]. This should obviously be done maintaining a certain level of security on the system. Additionally, it is important to state that implementing cryptography in such networks is done by means of processors that are not powerful enough to manage huge key sizes, mainly because the introduction of such systems should be somehow seamless to the existing communication infrastructure and cheap [^1] We can speculate that QKD will not be a major player to secure such networks due the fact that they are very costly as well as they require significant infrastructure to be deployed.. It is in this sense that, the inclusion of PQC in industrial and critical environments poses an interesting trade-off between the benchmarks of those protocols. As mentioned previously, protecting ICS and CI from possible cyberattacks is fundamental due to the immense impact that those systems have in society and industry, making their failure to cause intolerable economic losses and, in the worse scenario, injuries and even casualties. Interestingly, this systems have shown to be vulnerable in the recent times with several hacking proposals [28, 29, 30, 31]. Therefore, the necessity of transitioning the security of industrial and CI to post-quantum cryptography is central to keep all those systems secure against a possible quantum threat, as it has been recently noted by the United States Cybersecurity and Infrastructure Security Agency (CISA), the National Security Agency (NSA) and the NIST [32].
I-A Motivation and Related work
Due to the global security threat posed by the possibility of a fault-tolerant quantum computer, PQC is one of the most important topics in cryptography at the moment. Thus, there are many works pointing out the importance and the lack of cybersecurity in OT environments as well as surveys about PQC cryptosystems. Specifically, there are many theoretical and experimental references regarding PQC cryptosystems, such as [33, 34, 35, 36]. However, there is a gap in the literature regarding the merge of both problems. Since OT environments are a clear and critical target for cybercriminals, it is of the most importance to protect those scenarios from quantum attacks as well. In comparison with previous works about PQC cited before, this manuscript aims to provide a perspective on the problem of implementing PQC algorithms in CI and OT environments. In the literature there are many works regarding the importance of cybersecurity and cryptography in OT [27, 28, 31, 37, 38, 39, 40, 41, 42]. As PQC algorithms were traditionally conceived from the point of view of IT communications, there are some requirements in industrial environments that are usually not fulfilled by those. Therefore, we want to emphasize the necessity of more research in PQC algorithms from the point of view of OT communications and, at the same time, more test benches implementing PQC algorithms in industrial environments. This comes with the objective of assessing their reliability for OT communications while providing information for cryptography researchers in order to develop new PQC algorithms that are tailored to fulfill those requirements. Papers such as [43, 34, 36] provide a good introduction of the problem and a survey for NIST PQC algorithms, while [44] gives a good introduction to PQC cybersecurity in OT. In the context of industrial scenarios it is common to underestimate the importance of cybersecurity as well as to consider it as a toll to productivity due to the additional costs. Moreover, each global agent is aiming to standardise PQC protocols in an independent manner, requirements that will be necessary to all vendors to fulfill once established. The NIST standardisation process stands as one of the first efforts for PQC standardisation and, while followed by a considerable amount of occidental countries, it is not the only one in the world, where countries such as France [45] or Germany [46] are also following different processes. Thus, it is very unlikely that a global adoption of the same standard will happen for this new field of cryptography, as it happened for the widespread RSA and ECC cryptographic schemes.
I-B Outline and contribution
In this context, the principal objective of this contribution is to stress out the necessity of the integration of new generation PQC protocols to industrial and CI environments as well as to discuss the state of affairs and challenges regarding such integration. Specifically, we aim to:
  • Provide an introduction to traditional cryptography in OT environments for industry experts for providing them the basic knowledge of the problem. As mentioned before, many industrial players may be unaware of the importance of integrating cryptography in their environments and, thus, it is our intention to provide them with the basic concepts. Also, this serves to introduce the challenges and requirements of integrating cybersecurity, in general, to CI and industrial environments.
  • Discuss how quantum computing can pose a threat to OT communications and show how this risk fundamentally differs from the one that IT communications may experience. This is aimed to show industrial players why transitioning to quantum secure cryptography will be critical as well as to show PQC developers how those networks should be protected.
  • Describe the state-of-the-art PQC families and protocols being considered, not only within the NIST standardization process but also within other processes around the world. This serves as an introduction of PQC cryptography for newcomers. Additionally, we intend to show that due to the uncertainty regarding the PQC algorithms that will be implemented (some strong candidates are being questioned or have been recently broken) and the fact that it seems that many global agents will adopt their own methods, PQC integration in OT environments will require a great degree of flexibility regarding implementation.
  • Provide a discussion of the state-of-affairs regarding PQC implementation in OT environments. Specifically, we want to pose the main challenges when integrating those quantum secure protocols in such scenarios. Finally, to show the necessity of more active research on this topic both by cryptographers and industrial players.
The manuscript is organized as follows: in Section II we provide a review of the communication systems in industrial environments as well as of the stringent requirements for integrating cybersecurity in them. We follow, in Section III, with a short review on cryptography and the threat that quantum computers pose to traditional ciphering schemes. Existing proposals of PQC algorithms are surveyed in Section IV presenting several protocols proposed by many worldwide agents. In addition, the performance benchmarks of those PQC candidates are presented. An overview of the state of PQC in industrial and CI environments is finally presented in Section V, were we speculatively discuss which existing protocols may be the ones for integration in those scenarios.
I-C Review Methodology
The methodology rearding the literature review conducted was as follows. An exhaustive online search was conducted in order to identify key works reviewing classical cryptography and PQC algorithms. For such initial search, the traditional databases for classical cryptography were explored, i.e. IEEE Xplore, ACM and the Cryptology ePrint Archive. Many references regarding PQC were found in those sources, but the review by Bernstein and Lange was specially useful to identify many lines and references regarding such field [21]. Furthermore, the NIST standardization process was also an starting point to identify many of the methodologies that present the potential to be implemented. Since the NIST standardization stands as the mainstream process in this line, it was the starting point to identify PQC protocols that go beyond the basic theory for PQC families. Once studied, we followed to other standardization efforts since one of the points of this perspective is to show that many global agents are independently doing such process and we wanted to show the heterogeneous nature of this field. Moreover, (post-quantum) cryptogrpahy is a dynamic field so many related blogs, e.g. Cloudflare [^2] https://blog.cloudflare.com/ or Google Bughunters [^3] https://bughunters.google.com/blog, were regularly read in order to follow developments of those fields at the time of writing this article. Regarding industrial cryptography, we based on the knowledge of one of the authors (G. Vidal) in order to identify a preliminary batch of relevant literature regarding this topic. Afterwards, we complemented such literature by means of IEEE Xplore and ACM databases as well as by getting references from such initial set of articles. Last but not least, some missing references were pointed by the referees, which were discussed for completeness of the work.
Once the literature was collected and understood, all the information was used to make discussions regarding the implementation of PQC in industrial networks as well as to identify which are the challenges associated to it. This has been done by comparing the core problem of IT cryptography in contrast to OT cryptography, which is the one tackled in this perspective. In this way, it has been seen that most of the PQC protocols that have been proposed would have difficulties to be integrated in industrial networks as they have been constructed from the IT point of view. Finally, future work has been pointed out in order to make such scenarios quantum secure.
@@ II Industrial environments
ICS are the components of the industrial sector and infrastructures, from essential services such as energy, water, transportation systems to manufacturing plants, agricultural systems, building automation systems, etc. In these infrastructures we will always find complex components that share a common denominator: physical processes that are modified by logical computation or viceversa. These components are called Cyber-Physical Systems (CPS). In Fig. 1, we show an example of ICS. In this example, in short, the instrumentation sensors measure physical variables, programmable logic controllers (PLCs) implement a control loop and send signals to the actuators. All the process is controlled and monitored by the Supervisory Control And Data Acquisition (SCADA) and the operator interacts via the Human-Machine Interface (HMI) or the Engineering Work Station (EWS).
@@
Since these systems control real world processes, any potential cyberattack impact on them could imply a physical effect in the real world. Hence, cyber-risks could turn not only into production downtimes but also physical damage to operators or users. A good survey of different objectives and techniques used to attack ICS networks has been collected in MITRE ATT&CK  [^4] https://attack.mitre.org/.
ICS and CPS are a part of larger infrastructures which interact with information technology (IT) systems at certain point. In Fig. 2, we show how IT and operational technology (OT) interact in this type of infrastructures according to the Purdue model, even though there are several models such as RIA 4.0 and others.
@@ II-A Differences between IT and OT
In summary, the cost of achieving a high level of cybersecurity can be much higher in older industrial plants due to the need for retrofitting and upgrading legacy systems. In contrast, new companies and modern facilities have the advantage of incorporating cybersecurity measures at a lower cost from the outset, due to the availability of cyber-embedded components and systems. However, it is crucial for all organizations, regardless of age, to prioritize cybersecurity to protect critical infrastructure and assets from evolving cyber threats.
II-B Industrial Cybersecurity Standards and Mechanisms
@@
In addition to broader standards and regulations, there are specific standards for components and systems used in industrial environments. These may include industrial communication protocols and HW requirements. By adhering to these standards, organizations can ensure compatibility and security among different components and systems within their infrastructure. The interoperability provided by these specific standards is crucial for maintaining a secure and efficient industrial ecosystem. Fig. 3 summarizes these interdependencies.
@@ III-A Overview on cryptography
In the ensuing paragraphs, we provide a comprehensive explanation of both schemes based on a network composed of three parties: a server, a user who wants to have a secure connection with it and an eavesdropper who wants to obtain information about the message. In cryptography those are usually referred to as Alice, Bob and Eve, respectively.
III-A1 Symmetric Key Schemes
@@
III-A2 Asymmetric Key Schemes
As defined above, asymmetric key or public key schemes are defined as cryptosystems where each of the parties involved in the communication use their own key to secure the information, the private key and the public key. Each of those are usually employed to share a secret between two parts and, therefore, this encryption schemes may be used for the confidential key exchange required in symmetric schemes, called Key Encapsulation Mechanism (KEM). In this sense, this could refer to the transmission of the bit string used as the input of an algorithm which can be used by Alice and Bob to generate the same symmetric key. Otherwise, if it is used to encrypt the message it is called Public Key Encapsulation (PKE). In asymmetric key schemes confidentiality, integrity and authenticity are achieved in the following ways: